Inregrate "\int \frac{\cos x dx}{\sin^2 x}"


1
Expert's answer
2022-05-31T06:56:11-0400

"\int \cfrac{\cos x dx}{\sin^2x} = [t = \sin x, dt = \cos x dx] = \int \cfrac{dt}{t^2} = -\cfrac{1}{t}+C=-\cfrac{1}{\sin x}+C"